A flask of 7L contains a mixture of 6g of O2 and 15g of N2 at 30°C.What is the partial pressure of the oxygen gas?

[1] 0.67 atm

[2] 1.9 atm

[3] 2.59 atm

[4] 22.1 atm

[5] None of the above


1
Expert's answer
2022-06-14T08:43:04-0400

V = 7 L

T = 30°C = 303 K

n(O2) = 6 / 32 = 0.1875 mol

n(N2) = 15 / 28 = 0.536 mol

P = ?


"P=~\cfrac {nRT}{V}=~\cfrac{ 0.1875*8.31*303}{7}= ~67.445 ~ kPa = ~0.67 ~atm"



Answer: [1] 0.67 atm